Calculate the number of moles of aluminum, sulfur, and oxygen atoms in 7.00 moles of aluminum sulfate, Al2(SO4)3
1
Expert's answer
2019-05-23T07:59:26-0400

2moles(Al)*7moles=14moles aluminum

3moles(S)*7moles=21moles sulfur

12moles(O)*7moles=84 moles oxygen